oguinns:

A population starts with 3000 and has a doubling time of 5 years. Write a growth or decay formula for the exponential function.

Vocaloid:

meant to get to this earlier but had trouble thinking of a way to explain w/o giving it away ;_; anyway, the general formula for exponential growth is P = A(1+r)^t typically where A = the initial population, r is the rate of increase, t is time we can modify this equation to account for the doubling time of five years by dividing the t by 5 and changing 1 + r to 2 because this is a doubling equation P = A(2)^t/5 then plug in the initial population and doubling time